# VeOmni: Scaling Any Modality Model Training with Model-Centric Distributed Recipe Zoo

### v1

@@ 1 Introduction
Extending existing training frameworks to omni-modal LLMs is non-trivial. To address the growing parameter sizes, modern training frameworks leverage various distributed training strategies [51, 40, 32, 41] to alleviate computation and memory bottlenecks. For example, Megatron-LM [51, 41] provides optimized transformer blocks [61] with advanced parallelization strategies like tensor parallelism (TP) and pipeline parallelism (PP). However, as omni-modal architectures become complex in both functionality and parameter size, directly applying these techniques to omni-modal LLMs often leads to load imbalance and poor scalability [17], due to the current frameworks tightly couple model definition with parallel logic. This entanglement hiders generalization to more diverse model architectures. Although recent efforts [24, 77, 17] attempt to mitigate the inefficiencies of omni-modal training, they still suffer from the coupling of communication and computation, resulting in significant engineering cost and limited extensibility.
@@ 2.2 LLM Training Frameworks
The landscape of large language model training frameworks has evolved significantly to address the computational demands of scaling. For pure text training, several mature frameworks have established themselves as industry standards. Megatron-LM [51, 41] pioneered optimized transformer blocks with advanced parallelization strategies including tensor parallelism (TP) and pipeline parallelism (PP), becoming the foundation for many subsequent frameworks. Colossal-AI [33] extends this paradigm by offering comprehensive 3D parallel training strategies that combine data, tensor, and pipeline parallelism for enhanced scalability. NeMo [50] provides an end-to-end cloud solution specifically designed for training large-scale LLMs with enterprise-grade features, while TorchTitan [36] and veScale [62] represent the newer generation of PyTorch-native frameworks that emphasize auto-parallelization and simplified programming models.
@@ B.1 OmniData Process and Collator API
from omniscale.data import (
@@ B.2 Building Omni-Modal LLMs
from omniscale.models import build_omni_model
@@ B.4 Parallel State of VeOmni
Listing LABEL:lst:veomni-parallel-state demonstrates the core interface of VeOmni’s parallel_state API, which provides a unified and declarative initialization of all N-dimensional parallel configurations. Unlike traditional process group–based parallel management, VeOmni leverages a DeviceMesh abstraction to represent and organize various parallel topologies, including data parallel (DP), tensor parallel (TP), expert parallel (EP), pipeline parallel (PP), and Ulysses-based sequence parallel (SP). Once initialized via init_parallel_state, users can conveniently access mesh and group objects, ranks, and communication topologies through the global parallel_state object. This design simplifies the orchestration of hybrid-parallel strategies and decouples low-level distributed logic from model implementation.
@@
from omniscale.distributed.parallel_state import get_parallel_state, init_parallel_state
@@ B.5 FSDP and parallelized
from omniscale.distributed.torch_parallelize import build_parallelize_model
@@ B.6 Ulysess flash attention forward Function
from omniscale.distributed.parallel_state import get_parallel_state
from omniscale.distributed.sequence_parallel import (
@@ B.7 Expert Parallel Plan and Expert Parallel Function
from omniscale.distributed.parallel_plan import ParallelPlan

### v2

@@ 1 Introduction
Extending existing training frameworks to omni-modal LLMs is non-trivial. To address the growing parameter sizes, modern training frameworks leverage various distributed training strategies [51, 40, 32, 41] to alleviate computation and memory bottlenecks. For example, Megatron-LM [51, 41] provides optimized transformer blocks [61] with advanced parallelization strategies like tensor parallelism (TP) and pipeline parallelism ($P$). However, as omni-modal architectures become complex in both functionality and parameter size, directly applying these techniques to omni-modal LLMs often leads to load imbalance and poor scalability [17], due to the current frameworks tightly couple model definition with parallel logic. This entanglement hiders generalization to more diverse model architectures. Although recent efforts [24, 77, 17] attempt to mitigate the inefficiencies of omni-modal training, they still suffer from the coupling of communication and computation, resulting in significant engineering cost and limited extensibility.
@@ 2.2 LLM Training Frameworks
The landscape of large language model training frameworks has evolved significantly to address the computational demands of scaling. For pure text training, several mature frameworks have established themselves as industry standards. Megatron-LM [51, 41] pioneered optimized transformer blocks with advanced parallelization strategies including tensor parallelism (TP) and pipeline parallelism ($P$), becoming the foundation for many subsequent frameworks. Colossal-AI [33] extends this paradigm by offering comprehensive 3D parallel training strategies that combine data, tensor, and pipeline parallelism for enhanced scalability. NeMo [50] provides an end-to-end cloud solution specifically designed for training large-scale LLMs with enterprise-grade features, while TorchTitan [36] and veScale [62] represent the newer generation of PyTorch-native frameworks that emphasize auto-parallelization and simplified programming models.
@@ B.1 OmniData Process and Collator API
from veomni.data import (
@@ B.2 Building Omni-Modal LLMs
from veomni.models import build_omni_model
@@ B.4 Parallel State of VeOmni
Listing LABEL:lst:veomni-parallel-state demonstrates the core interface of VeOmni’s parallel_state API, which provides a unified and declarative initialization of all N-dimensional parallel configurations. Unlike traditional process group–based parallel management, VeOmni leverages a DeviceMesh abstraction to represent and organize various parallel topologies, including data parallel (DP), tensor parallel (TP), expert parallel (EP), pipeline parallel ($P$), and Ulysses-based sequence parallel (SP). Once initialized via init_parallel_state, users can conveniently access mesh and group objects, ranks, and communication topologies through the global parallel_state object. This design simplifies the orchestration of hybrid-parallel strategies and decouples low-level distributed logic from model implementation.
@@
from veomni.distributed.parallel_state import get_parallel_state, init_parallel_state
@@ B.5 FSDP and parallelized
from veomni.distributed.torch_parallelize import build_parallelize_model
@@ B.6 Ulysess flash attention forward Function
from veomni.distributed.parallel_state import get_parallel_state
from V e Omni.distributed.sequence_parallel import (
@@ B.7 Expert Parallel Plan and Expert Parallel Function
from veomni.distributed.parallel_plan import ParallelPlan